OpenStudy (anonymous):

what is the 25th term of the arithmetic sequence 2,1 3/5, 1 1/5??

OpenStudy (anonymous):

If you consider the 2 to be the first term, then it is -38/5.

OpenStudy (anonymous):

If you consider 2 to be the zeroth term, then it is -8.

OpenStudy (anonymous):

I think it is the first term

OpenStudy (amistre64):

if we subtract 2 from all terms we get it into a set up that we can see how much is add to each term

OpenStudy (amistre64):

2 ,1 3/5, 1 1/5 -2 -2 -2 -------------- 0, -2/5 , .....

hero (hero):

amistre, How come you haven't given me any of your medals yet?

OpenStudy (amistre64):

oh but i have :)

OpenStudy (amistre64):

an = 2 -2/5(n-1) is the set up for the formula then a25 = 2 -2/5 (24) a25 = 2 -48/5 a25 = 10/5 -48/5 = -38/5